OpenStudy (anonymous):

will give metal to best answer - simplify the product. -4 times 10 times 6 A. -240 B. -241 C. -238 D. -40

OpenStudy (anonymous):

A

OpenStudy (anonymous):

For this one there isn't gonna be a detailed explanation

OpenStudy (anonymous):

lol, ok

OpenStudy (anonymous):

It's just done on a calculator

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

A because -4*10=-40*6=-240